Find the area bounded by y = (11 - x)^3, the lines y=x and the x-axis. *


1
Expert's answer
2021-10-28T07:17:52-0400

Explanation:

According to second theorem of Guldino, volume obtined by a rotation of a section bounded by a function f(x) 

and the x-axis between a and b is :"V=\pi\int_a^bf^2(x)dx"


In our case, we have 

"V= \pi \int_1^2e^{(2x)}dx=\pi/2(e^4-e^1)"

so it will be: "Y=(11-x)^3"

"=(11-4)^3=343"